OpenStudy (anonymous):

solve: 1/81=b^-4 Can you please show me the steps to take? Thanks

OpenStudy (anonymous):

you could start with \[81=b^4\] then \[b=\sqrt[4]{81}\]

OpenStudy (anonymous):

I'm sorry i don't understand.

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

\[\frac{ 1 }{ 81 }=b ^{-4},\frac{ 1 }{ 3^4 }=b ^{-4},3^{-4}=b ^{-4},b=?\]